OpenStudy (anonymous):

Need help with capacitance: Hey guys , so I've been reviewing the question and I really got stuck at some of the questions , I'll post them with my solution and you guys please help me if I was wrong: *Note some of the question I could not solve them..

OpenStudy (anonymous):

Q1/An isolated charged conducting sphere of radius 12 cm creates an electric field of 4.90 x 10^4 N/C at a distance 21 cm from its center. (a) What is its surface charge density? (b) What is its capacitance? a) Surface q density = q/a and we want q so we use e=kq/r^2 Q=9*9^10/4.90*10^4*(21*10^-3)^2 after finding q we find the surface q denisity which is: =Q/4*(3.14*(12*10^-3))^2 is it right? b) c=q/delta v we find delta v which is v=ed which is 4.90*10^4*(21*10^-3) then we use it to find c. is my solutions right ? I'll post the other questions..

OpenStudy (anonymous):

Q2/An air-filled capacitor consists of two parallel plates, each with an area of 7.6 cm2, separated by a distance of 1.80 mm. A 20 V potential difference is applied to these plates. Calculate (a) the electric field between the plates, (b) the surface charge density, (c) the capacitance, and (d) the charge on each plate. a) we use v=ed so e=v/d so e=20/0.0018000 we will get e. b) it's = to q/a and we find q by: q=k/e*r^2 like we did in question 1 and then we use it to find the surface charge which is q/a so our q/0.00076m^2 c) c=q/delva v we use our q then divide it by 20 d) I don't know how to solve this :(

OpenStudy (anonymous):

Q3/When a potential difference of 150 V is applied to the plates of a parallel-plate capacitor, the plates carry a surface charge density of 30 nC/cm2. What is the spacing between the plates? I don't know how to solve this really :( ------ Q4/ Two capacitors, C1= 5 µ F and C2 =12 µ F, are connected in parallel, and the resulting combination is connected to a 9 V battery. (a) What is the equivalent capacitance of the combination? What are (b) the potential difference across each capacitor and (c) the charge stored on each capacitor? a) it's easy no need.. b) how could I find the potential difference here ? v=ed ?? c) it's the same as the second question that I couldn't solve :(

ganeshie8 (ganeshie8):

Q1/An isolated charged conducting sphere of radius 12 cm creates an electric field of 4.90 x 10^4 N/C at a distance 21 cm from its center. (a) What is its surface charge density? (b) What is its capacitance? a) Surface q density = q/a and we want q so we use e=kq/r^2 Q=9*9^10/4.90*10^4*(21*10^-3)^2 after finding q we find the surface q denisity which is: =Q/4*(3.14*(12*10^-3))^2 is it right?

ganeshie8 (ganeshie8):

e=kq/r^2 \(q = er^2/k\) \(e = 4.90 x 10^4 \) \(k = 9*10^9\) \(r = 21*10^-2\)

OpenStudy (anonymous):

oh , thanks , I got it wrong :(

ganeshie8 (ganeshie8):

np :)

ganeshie8 (ganeshie8):

u want me check others ?

OpenStudy (anonymous):

Yeah please :(

ganeshie8 (ganeshie8):

okay... 100cm = 1m so, 1cm = 10^-2 m ok

ganeshie8 (ganeshie8):

it seems u were using 1cm = 10^-3 m in all ur calculations

OpenStudy (anonymous):

oh boy :( it's ok do not worry about that , the only thing I'm worried about is solving the problem itself , is my solution ok ? I really hate this subject and thank god it's the last :)

ganeshie8 (ganeshie8):

Q2/An air-filled capacitor consists of two parallel plates, each with an area of 7.6 cm2, separated by a distance of 1.80 mm. A 20 V potential difference is applied to these plates. Calculate (a) the electric field between the plates, (b) the surface charge density, (c) the capacitance, and (d) the charge on each plate.

ganeshie8 (ganeshie8):

d) I don't know how to solve this :(

ganeshie8 (ganeshie8):

Multiply below two quantities :- 1) Area of plate 2) Answer from part b

OpenStudy (anonymous):

why did you jump to d ? is my solution for the others correct ? thanks buddy

ganeshie8 (ganeshie8):

it looks okay, \(\epsilon_r = 1\) for air right ?

OpenStudy (anonymous):

I don't really know as I said I hate this subject I'm not that good when it comes to this subject :(

ganeshie8 (ganeshie8):

its okay, you're just doing the dry electric field stuff... the interesting magnetic fields/motors/generators starts soon, then u wil start to like the subject more

OpenStudy (anonymous):

I'm afraid this is the last time I'll study this subject , since I'm majoring in computer science :) anyways what is the solution for d ? and how about the other questions ?

ganeshie8 (ganeshie8):

you can work d) in atleast two ways :- \(Q = CV\)

ganeshie8 (ganeshie8):

the other way, i have told u already

ganeshie8 (ganeshie8):

\(Q = CV\) V = 20 C = Answer from part c

OpenStudy (anonymous):

ah thanks , how about the other questions ?

OpenStudy (anonymous):

I just wanted to say that I appreciate your effort to solve this , and helping me understand , thanks a lot :)

ganeshie8 (ganeshie8):

3rd q :-

OpenStudy (anonymous):

oh thank you so so so much , you helped a lot :$

ganeshie8 (ganeshie8):